OpenStudy (anonymous):

C=5/p(F-32) ...Solve for "F"

OpenStudy (anonymous):

P=9

mathslover (mathslover):

given : \[\large{C = \frac{5}{p(F-32)}}\] given P = 9 so : \[\large{\frac{5}{9(F-32)}}\]

mathslover (mathslover):

m i right? @gabbbdee

OpenStudy (anonymous):

Yes the bottom one is the correct equation.

OpenStudy (anonymous):

wait... no

OpenStudy (anonymous):

its C=5/9 (F-32) and i have to solve for F

mathslover (mathslover):

ok so it is : \[\large{C=\frac{5}{9}(F-32)}\]

mathslover (mathslover):

right?

OpenStudy (anonymous):

Do I multiply "(F-32)" by 5/9 first?

OpenStudy (anonymous):

yes.

mathslover (mathslover):

No ... first step will be : \[\large{C\times \frac{9}{5}=\frac{5}{9}(F-32) \times \frac{9}{5}}\]

mathslover (mathslover):

I just multiplied both sides by 9/5

OpenStudy (anonymous):

why do you flip the fraction?

mathslover (mathslover):

just to cancel 5/9 there

mathslover (mathslover):

just for example: \[\large{3=\frac{a}{3}}\] \[\large{3\times 3 = \frac{a}{3} \times 3}\] \[\large{9 = \frac{a}{\cancel{3}^1}\times \cancel{3}^1}\] \[\large{9=a}\]

mathslover (mathslover):

that above was an example

OpenStudy (anonymous):

ohhh okay, and after i muliply both sides whats next? can you go through the problem and explain it?

mathslover (mathslover):

ok \[\large{C\times \frac{9}{5} = \frac{5}{9} \times \frac{9}{5} (F-32)}\] \[\large{C\times \frac{9}{5} = \frac{\cancel{5}^1}{\cancel{9}^1}\times \frac{\cancel{9}^1}{\cancel{5}^1} (F-32)}\] \[\large{\frac{9C}{5}=F-32}\]

mathslover (mathslover):

\[\large{9C=(F-32)5}\] can you expand 5(F-32) ?

mathslover (mathslover):

@gabbbdee please reply

OpenStudy (anonymous):

sorry if i expand it be 5F-160 of i do the rainbow thing and the equation will equal -32 if im right? O.o

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

\[\large{9C=5F-160}\] \[\large{9C+160=5F}\] \[\large{\frac{9C+160}{5}=F}\]

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

@nincompoop lol